Find the sum of product expansion of the Boolean function f(x,y,z) = (x+z)y


1
Expert's answer
2020-12-30T15:42:06-0500

"F(x,y,z)=(x+z)y"

To determine the sum of product expansion of "F", we could use boolean identities or construct a table that determines all values of "F". I will be using boolean identities.


"F(x,y,z)=(x+z)y"

"=xy+zy \text{ (Distributive Law) }\\
=xy.1+zy.1 \text{ (Identity Law) }\\
=xy(z+ \bar{z})+zy(x+\bar{x}) \text{ (Unit Property) }\\
=xyz+xy\bar{z}+zyx+zy\bar{x} \text{(Distributive Law) }\\
=xyz+xy\bar{z}+xyz+\bar{x}yz \text{ (Commutative Law) }\\
=xyz+xy\bar{z}+\bar{x}yz \text{(Idempotent Law) }"